Circular No. 8220

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

CXOU J010043.1–721134 R. C. Lamb and T. A. Prince, California Institute of Technology; D. J.

Macomb, Boise State University; and W. A. Majid, Jet Propulsion Laboratory, report that the possible anomalous x-ray pulsar (AXP) designated CXOU J010043.1–721134 has a pulse period of 8.02 s, rather than the previously reported 5.4-s period (Lamb et al. 2002, Ap.J. 574, L29) from Chandra ACIS-I data, which yield a Nyquist frequency of 0.154 Hz (6.48s period): “In the Fourier transform of these data, two strong peaks at 0.125 and 0.184 Hz were observed with power values of 40, where the average power is normalized to 1. ROSAT data for this same source favored the higher frequency. However, new data for this source, from the XMM-Newton satellite (observation 0018540101), clearly show that the correct frequency is near 0.125 Hz. The XMM data were obtained with the PN detector, which has a digitization interval of 0.0734 s, giving a Nyquist frequency of 6.8 Hz, well above the region of interest. The times of arrival of 4048 source photons, collected in the 27300-s observation (epoch 2001 Nov. 21.015 UTC), were corrected to the solar-system barycenter. The Fourier transform of these corrected times show a strong peak at 0.124699(5) Hz with a normalized power value of 28. Power peaks at the second and third harmonics are present as well. The luminosity of the source (0.2–10 keV), corrected to the 60000-pc distance of the Small Magellanic Cloud, is 2.6 × 10^{35} erg/s. By combining this observation with the Chandra observation made 189.934 days earlier, we derive a frequency time derivative of −4 ± 3 × 10^{−13} s. The relative constancy of the luminosity of this source, its magnitude, and its small spin-down rate (time scale P/Ṗ ∼ 10000 yr) are consistent with the characteristics of AXPs (cf. Mereghetti et al. 2002,

http://xxx.lanl.gov/abs/astro-ph/?0205122).”

2003 SS_{84}

M. C. Nolan, A. A. Hine, and E. S. Howell, National Astronomy and Ionosphere Center; and L. A. M. Benner and J. D. Giorgini, Jet Propulsion Laboratory, report that Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm) obtained on Oct. 4–8 show that 2003 SS_{84} (cf. MPEC 2003-S40) is a binary system. Preliminary estimates of average diameters, based on range extents at 15-m resolution, are 120 and 60 m. The pair was seen at approximately the same orbital phase at Oct. 4.073, 5.083, 7.034, and 8.003 UT, indicating an orbital period of just under 24 hr or an integer fraction thereof. Optical lightcurve observations of this object would be valuable.

2003 October 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green